# Vision Expert-guided Visual Tokenizer and Manufacturing Process

### v1

@@ 4.2 Ablations
Effect of InstructIAD. InstructIAD is placed as an instruction-tuning dataset for IAD, that aims to align the visual features and textual descriptions, leading to a performance gain compared with the base model (LLaVA-1.6-mistral-7B). Notably, as shown in Tab. 3, the performance with manufacturing processes also increases, empirically demonstrating the intrinsic relationship between attribute-level recognition and manufacturing-aware defect analysis.
@@
Effect of EG-RoI. Comparative analysis in Tab. 3 (row 3 vs. 5) confirms EG-RoI’s consistent performance gains across both manufacturing-aware and conventional detection scenarios. We further evaluate vision expert effectiveness through size-based defect categorization on MVTec-AD, adopting MS-COCO’s partitioning protocol [21]: small ($<0.01$ image area), medium ($0.01\sim 0.1$), and large ($>0.1$). As shown in Tab. 6, EG-RoI significantly outperforms LLaVA-1.6’s AnyRes module on small/medium defects. The baseline exhibits severe classification bias with 100% defect detection rate but 16.1% false positives on normal samples. Table 5 reveals two critical findings: (1) Anomaly map quality directly impacts detection accuracy, validating our expert integration strategy; (2) Triad achieves competitive performance even without annotated regions (AnyRes mode), surpassing conventional methods. Notably, expert modules maintain manufacturing-aware improvements despite random box noise injection (Null configuration).
Ablations on the different manufacturing process. For systematic evaluation, we collect three distinct manufacturing process variants for MVTec AD [2] through multi-source construction: (1) Internet: Domain-specific processes collected from internet for all 15 product categories; (2) LLM Generation: Automated process synthesis using general-purpose LLMs (LLaMA-3 [9] and GPT-4 [1]). As evidenced in Tab. 4, Triad demonstrates consistent performance gains across all process variants (0.7%-2.5% accuracy improved), empirically validating its dual capability in manufacturing process comprehension and generalization to unseen industrial workflows. Extended process examples are provided in Sec. B.
Table 3: Ablations on finetuning with different components: InstructIAD, CoT-M, and EG-RoI tokenizer. The base model is LLaVA-1.6-mistral-7B [22]
@@
76.9% 75.9% (1.0%) $\downarrow$
78.8% 80.4% (1.6%) $\uparrow$
79.8% 82.5% (2.7%) $\uparrow$
85.4% 83.9% (1.5%) $\downarrow$
85.0% 87.5% (2.5%) $\uparrow$
Table 4: Ablation on the robustness of the manufacturing process on MVTec-AD [2] Three types of the manufacturing processes of 15 products on MVTec-AD are collected from the Internet (Web), generated by llama3 [9] (LLM), and generated by GPT4 [1] (GPT).
@@
Qwen2-VL [31] 2B 77.0% 51.7%(25.3%) $\downarrow$ 46.7%(30.3%) $\downarrow$ 46.8%(30.2%) $\downarrow$
LLava-1.6 [22] 7B 76.9% 77.6% (0.7%) $\uparrow$ 75.9% (1.0%) $\downarrow$ 77.3% (0.4%) $\uparrow$
LLaVA-OneVision-si [16] 7B 77.7% 69.4% (8.3%) $\downarrow$ 60.6%(17.1%) $\downarrow$ 67.2%(10.5%) $\downarrow$
LLaVA-OneVision-ov [16] 7B 91.0% 88.8% (2.2%) $\downarrow$ 80.8%(10.2%) $\downarrow$ 87.1% (3.9%) $\downarrow$
Qwen2-VL [31] 7B 84.4% 70.9%(13.5%) $\downarrow$ 61.1%(23.3%) $\downarrow$ 67.1%(17.3%) $\downarrow$
Qwen2-VL [31] 72B 87.1% 85.2% (1.9%) $\downarrow$ 79.5% (7.6%) $\downarrow$ 82.6% (4.5%) $\downarrow$
Myriad [19] 7B 79.3% 78.5% (0.8%) $\downarrow$ 81.5% (2.2%) $\uparrow$ 81.5% (2.2%) $\uparrow$
Triad-llava-1.6 7B 85.0% 85.7% (0.7%) $\uparrow$ 87.5% (2.5%) $\uparrow$ 86.4% (1.4%) $\uparrow$
Triad-ov 7B 91.2% 91.8% (0.6%) $\uparrow$ 92.6% (1.4%) $\uparrow$ 92.2% (1.0%) $\uparrow$
@@
Null - 83.3% 83.4% (0.1%) $\uparrow$
AnyRes - 84.0% 85.1% (1.1%) $\uparrow$
April-GAN [7] 87.6% 83.0% 85.0% (2.0%) $\uparrow$
AnomalyClip [40] 91.1% 84.6% 86.1% (1.5%) $\uparrow$
MuSc [18] 97.3% 85.0% 87.5% (2.5%) $\uparrow$
@@ A.1 Results on PCB-Bank
We also test results on PCB-Bank [33], which includes a bunch of PCB images in a real industrial scene. Because part of the images in PCB-Bank come from VISA [41] and RealIAD [30], which are overlapped with our training data, we use the rest subset of PCB-Bank to evaluate for maintaining the zero-shot setting. Detecting defects on PCB is complicated and needs a lot of experience and knowledge in electronics, which is lacking in the general LMMs’ knowledge base. However, Triad is still outperformed due to its ability to use extra knowledge from the detailed manufacturing processes of PCB products.
@@
Qwen2-VL [31] 2B 49.9% 61.0% (11.1%) $\uparrow$
LLava-1.6 [22] 7B 38.9% 59.9% (21.0%) $\uparrow$
MiniCPM-V [34] 8B 61.2% 61.0% (0.2%) $\downarrow$
LLaVA-OneVision-si [16] 8B 60.8% 61.0% (0.2%) $\uparrow$
LLaVA-OneVision-ov [16] 8B 59.6% 60.9% (1.1%) $\uparrow$
Qwen2-VL [31] 7B 59.5% 61.1% (1.6%) $\uparrow$
Qwen2-VL [31] 72B 59.6% 61.0% (1.4%) $\uparrow$
LLaVA-OneVision-ov [16] 72B 60.9% 61.4% (0.5%) $\uparrow$
Myriad [19] 7B 61.8% 61.8% (0.0%) $\uparrow$
Triad-llava-1.6 7B 63.6% 64.8% (1.2%) $\uparrow$
Triad-ov 8B 62.8% 63.8% (1.0%) $\uparrow$
A.2 Extensive qualitative results
@@
Another example from PCB-Bank [33] is present in Fig. A2. The models are tested in two settings: zero-shot anomaly detection (green box) and anomaly detection with manufacturing process information (pink box). In the zero-shot anomaly detection setting, all models fail to correctly identify the missing button, although Triad detects a missing substrate that might be part of the design. This indicates a need for anomaly detection that incorporates information from the manufacturing process. The anomaly detection with manufacturing process information is presented on the pink boxes. In this example, Triad demonstrates its superior ability to utilize information from the PCB manufacturing process and precisely identify the missing button.
@@ B More details for evaluation
Evaluation prompts are following templates shown in Tab. A5. For general LMMs and our method, questions and options are provided with additional instruction leading to a binary output of “A” or “B” for the convenience of calculating accuracy. For the LMM-based IAD methods, we keep their original prompts as much as possible. Specifically for AnomalyGPT [10], it has a hint before the question on MVTec AD [2].
@@ C Task details
We have manually annotated detailed IAD captions on VisA [41] and a part of RealIAD [30]. An example is shown in Fig. A3 with a human-annotated caption. When labeling the IAD image, we first describe the product in the image and its detailed features following the general image-captioning task. Then, the defect is described by its type, location, texture, color, and shape. For multiple defects, each defect will be described separately with their relative position to the image. Later, captions are processed by Large Language Models (LLMs) to construct different task-based instruction datasets.
@@ D.1 Binarized anomaly segmentation map
Triad uses the binarized outputs of anomaly segmentation vision experts in EG-RoI. We observe that the range of anomaly scores varies between different vision experts, which makes it difficult to integrate them in a unified manner. The conventional normalization methods in  [7, 4, 12] are based on the entire test set, which is not feasible in real applications and requires a manually selected threshold for further binarization. Considering that vision experts only provide suspicious regions, we simply normalize the scores to the range 0-1 in one anomaly map and retain the scores that are larger than 0.9 on the anomaly map. As for threshold, Tab. A2 shows that although accuracy varies with different thresholds, Triad has the robustness against the increase of false anomaly regions (Pixel-FPR) and achieves a good performance when the True Positive Rate is low.
@@ D.2 Impact of the accuracy of anomaly maps
We calculate the accuracy of anomaly maps provided by vision experts for different thresholds on the MVTec AD test set in Tab. A3. Though selecting the best threshold, the performance of vision experts is typically poorer than Triad, and they also lack the ability to utilize extra context such as the manufacturing process to improve their performance on new products.
@@
$\Delta$ 2.5$\uparrow$ 1.0$\uparrow$ 1.2$\uparrow$ 1.3$\uparrow$ 1.8$\uparrow$ 0.7$\uparrow$ 1.3$\uparrow$ 0.8$\uparrow$ 0.8$\uparrow$ 0.5$\uparrow$ 0.2$\downarrow$
@@ E.1 Confidence Voting Mechanism
Confidence Voting Mechanism (CVM) is used in one-shot evaluation with the zero-shot model boosting the one-shot model in Traid (LLaVA-1.6) due to the lack of multi-image ability of LLaVA-1.6 [22]. In LMMs, references do not always bring benefits, instead, it could be misleading if the reference has some differences with the query image, Thus, the opinion of the zero-shot model is worth thinking if LMM is not strong enough for understanding multiple images. Confidence Voting Mechanism uses scores to collect the results of zero-shot model and one-shot model and give a combined result. As LMMs are bad at outputting confidence scores, the Confidence Voting Mechanism uses the option’s word predict score from the language model and compares them by following rules: If two models reach a consensus, directly give the result; else if one model gives a negative opinion, which is saying the query is non-defective, its word predict score would be compared with the reference predict score by the same model. If the model is trusting the query is more like a normal sample than the reference, it will be trusted and result in a non-defective judge, otherwise, the opposite opinion will be adopted. The reason for comparing with the reference predict score is that the space of predicted score by the two models is not the same and is unable to be normalized, so direct comparison makes no sense for the result. The effect of the Confidence Voting Mechanism is shown in Tab. A4.
model 1-shot base + MFG Proc.
0-shot model 85.0% 87.5% (2.5%) $\uparrow$
1-shot model 87.8% 88.3% (0.5%) $\uparrow$
1-shot model (CVM) 87.7% 88.4% (0.7%) $\uparrow$
@@ 4. Color Coding:
    - During the insulation process, the insulation is color-coded with different color to differentiate the wires within the cable.

### v2

@@ 4.2 Ablations
Effect of InstructIAD. InstructIAD is placed as an instruction-tuning dataset for IAD, which aims to align the visual features and textual descriptions, leading to a performance gain compared with the base model (LLaVA-1.6-mistral-7B). Notably, as shown in Tab. 3, the performance with manufacturing processes also increases, empirically demonstrating the intrinsic relationship between attribute-level recognition and manufacturing-aware defect analysis.
@@
Effect of EG-RoI. Comparative analysis in Tab. 3 (row 3 vs. 5) confirms EG-RoI’s consistent performance gains across both manufacturing-aware and conventional detection scenarios. We further evaluate vision expert effectiveness through size-based defect categorization on MVTec-AD, adopting MS-COCO’s partitioning protocol [21]: small (<0.01<0.01 image area), medium ($0.01\sim 0.1$), and large (>0.1>0.1). As shown in Tab. 6, EG-RoI significantly outperforms LLaVA-1.6’s AnyRes module on small/medium defects. The baseline exhibits severe classification bias with 100% defect detection rate but 16.1% false positives on normal samples. Table 5 reveals two critical findings: (1) Anomaly map quality directly impacts detection accuracy, validating our expert integration strategy; (2) Triad achieves competitive performance even without annotated regions (AnyRes mode), surpassing conventional methods. Notably, EG-RoI maintains manufacturing-aware improvements despite random box noise injection (Null mode).
Ablations on the different manufacturing process. For systematic evaluation, we collect three distinct manufacturing process variants for MVTec AD [2] through multi-source construction: (1) Internet: Domain-specific processes collected from the internet for all 15 product categories; (2) LLM Generation: Automated process synthesis using general-purpose LLMs (LLaMA-3 [9] and GPT-4 [1]). As evidenced in Tab. 4, Triad demonstrates consistent performance gains across all process variants (0.7%-2.5% accuracy improved), empirically validating its dual capability in manufacturing process comprehension and generalization to unseen industrial workflows. Extended process examples are provided in Sec. B.
Table 3: Ablations on fine-tuning with different components: InstructIAD, CoT-M, and EG-RoI tokenizer. The base model is LLaVA-1.6-mistral-7B [22]
@@
[…]
Table 4: Ablation on the robustness of the manufacturing process on MVTec-AD [2]. Three types of the manufacturing processes of 15 products on MVTec-AD are collected from the Internet (Web), generated by llama3 [9] (LLM), and generated by GPT4 [1] (GPT).
@@
[…]
@@
[…]
@@ 5 Conclusion
Acknowledgement
This work was supported by the National Key R&D Program of China under Grant No. 2023YFA1008500.
@@ A.1 Results on PCB-Bank
We also test results on PCB-Bank [33], which includes a bunch of PCB images in a real industrial scene. Because part of the images in PCB-Bank come from VISA [41] and RealIAD [30], which are overlapped with our training data, we use the rest subset of PCB-Bank to evaluate for maintaining the zero-shot setting. Detecting defects on PCB is complicated and needs a lot of experience and knowledge in electronics, which is lacking in the general LMMs’ knowledge base. However, Triad still outperforms due to its ability to use extra knowledge from the detailed manufacturing processes of PCB products.
@@
[…]
A.2 Results on MMAD
We report results on MMAD (categories for training are filtered out) in Tab. A2. Triad still achieves the best performance on MMAD and outperforms LLaVA-OV by 9% in both 0/1-shot. Note that the references for 1-shot evaluation on MMAD (especially for GoodsAD) are not actually the same product as the query image. Therefore, all models do not improve much under 1-shot setting.
Table A2: 0/1-shot anomaly detection performance on MMAD.
Model Qwen2-VL-7B Qwen2.5-VL-7B LLaVA-OV-ov Triad-ov
0-shot 56.25% 50.14% 60.74% 70.22%
[…]
1-shot 57.78% 42.71% 60.80% 70.92%
A.3 Extensive qualitative results
@@
Another example from PCB-Bank [33] is present in Fig. A2. The models are tested in two settings: zero-shot anomaly detection (green box) and anomaly detection with manufacturing process information (pink box). In the zero-shot anomaly detection setting, all models fail to correctly identify the missing button, although Triad detects a missing substrate that might be part of the design. This indicates a need for anomaly detection that incorporates information from the manufacturing process. The anomaly detection with manufacturing process information is presented in the pink boxes. In this example, Triad demonstrates its superior ability to utilize information from the PCB manufacturing process and precisely identify the missing button.
@@
A.4 Failure case study
Figure A4 presents four failure cases revealing Triad’s current limitations: (a) Viewpoint ambiguity: Requires 3D scans/X-ray imaging for better detection; (b) Design-related errors: Need CAD/GERBER design files for resolution; (c) Label-threshold conflict: Defects below a prescribed size are labeled as normal by the dataset; (d) Texture blindness: Complex patterns obscure flaws in textiles, which necessitates texture-aware training data.
@@ B More details for evaluation
Evaluation prompts are following templates shown in Tab. A6. For general LMMs and our method, questions and options are provided with additional instructions leading to a binary output of “A” or “B” for the convenience of calculating accuracy. For the LMM-based IAD methods, we keep their original prompts as much as possible. Specifically for AnomalyGPT [10], it has a hint before the question on MVTec AD [2].
@@ C Task details
We have manually annotated detailed IAD captions on VisA [41] and a part of RealIAD [30]. An example is shown in Fig. A3 with a human-annotated caption. When labeling the IAD image, we first describe the product in the image and its detailed features following the general image-captioning task. Then, the defect is described by its type, location, texture, color, and shape. For multiple defects, each defect will be described separately with its relative position to the image. Later, captions are processed by Large Language Models (LLMs) to construct different task-based instruction datasets.
@@
(a) Screw (MVTec-AD)
(b) PCB7 (PCB-Bank)
(c) Carpet (MVTec-AD)
(d) Grid cloth (WFDD)
Figure A4: Failure cases from MVTec-AD, WFDD, and PCB-Bank.
@@ D.1 Binarized anomaly segmentation map
Triad uses the binarized outputs of anomaly segmentation vision experts in EG-RoI. We observe that the range of anomaly scores varies between different vision experts, which makes it difficult to integrate them in a unified manner. The conventional normalization methods in  [7, 4, 12] are based on the entire test set, which is not feasible in real applications and requires a manually selected threshold for further binarization. Considering that vision experts only provide suspicious regions, we simply normalize the scores to the range 0-1 in one anomaly map and retain the scores that are larger than 0.9 on the anomaly map. As for threshold, Tab. A4 shows that although accuracy varies with different thresholds, Triad has robustness against the increase of false anomaly regions (Pixel-FPR) and achieves a good performance when the True Positive Rate is low.
@@ D.2 Impact of the accuracy of anomaly maps
We calculate the accuracy of anomaly maps provided by vision experts for different thresholds on the MVTec AD test set in Tab. A5. Though selecting the best threshold, the performance of vision experts is typically poorer than Triad, and they also lack the ability to utilize extra context, such as the manufacturing process to improve their performance on new products.
model 1-shot base + MFG Proc.
[…]
@@
[…]
@@ E.1 Confidence Voting Mechanism
Confidence Voting Mechanism (CVM) is used in one-shot evaluation with the zero-shot model boosting the one-shot model in Traid (LLaVA-1.6) due to the lack of multi-image ability of LLaVA-1.6 [22]. In LMMs, references do not always bring benefits; instead, they could be misleading if the reference has some differences from the query image. Thus, the opinion of the zero-shot model is worth thinking about whether LMM is not strong enough for understanding multiple images. Confidence Voting Mechanism uses scores to collect the results of zero-shot model and one-shot model and give a combined result. As LMMs are bad at outputting confidence scores, the Confidence Voting Mechanism uses the option’s word predict score from the language model and compares them by following rules: If two models reach a consensus, directly give the result; else if one model gives a negative opinion, which is saying the query is non-defective, its word predict score would be compared with the reference predict score by the same model. If the model is trusting the query more like a normal sample than the reference, it will be trusted and result in a non-defective judgment; otherwise, the opposite opinion will be adopted. The reason for comparing with the reference prediction score is that the space of predicted scores by the two models is not the same and unable to be normalized, so direct comparison makes no sense for the result. The effect of the Confidence Voting Mechanism is shown in Tab. A3.
@@ 4. Color Coding:
    - During the insulation process, the insulation is color-coded with different colors to differentiate the wires within the cable.